### v2

# Product Design Using a Generative Adversarial Network: Incorporating Consumer Preferences and External Data

Abstract
The development of generative artificial intelligence (AI) enables large-scale product design automation. However, this automated process usually does not incorporate consumer preference information from the internal dataset of a company. Furthermore, external sources such as social media and user-generated content (UGC) websites often contain rich product design and consumer preference information, but such information is not utilized by companies when generating designs. We propose a semi-supervised deep generative framework that integrates consumer preferences and external data into the product design process, allowing companies to generate consumer-preferred designs in a cost-effective and scalable way. We train a predictor model to learn consumer preferences and use predicted popularity levels as additional input labels to guide the training procedure of a continuous conditional generative adversarial network (CcGAN). The CcGAN can be instructed to generate new designs with a certain popularity level, enabling companies to efficiently create consumer-preferred designs and save resources by avoiding the development and testing of unpopular designs. The framework also incorporates existing product designs and consumer preference information from external sources, which is particularly helpful for small or start-up companies that have limited internal data and face the “cold-start” problem. We apply the proposed framework to a real business setting by helping a large self-aided photography chain in China design new photo templates. We show that our proposed model performs well in terms of generating appealing template designs for the company.
Keywords: Product design, generative models, deep learning, user-generated content (UGC)
1 Introduction
Product design is pivotal for firms to attract consumers and achieve success. The traditional product design process involves hiring professional human designers to produce new designs in small batches, which is a costly and unscalable approach. The rapid development of generative artificial intelligence (AI) can potentially make the task of generating new product aesthetic designs effortless. For instance, the big generative adversarial network (BigGAN) (Brock et al.,, 2018) is capable of generating high-resolution and diverse images across multiple classes. Platforms such as Artbreeder enable users to create unique artworks by blending existing images. [^1] See https://www.artbreeder.com/. From product packaging to car components and retail displays, industrial designers can use generative AI to explore more ideas and develop initial design concepts significantly faster than they could with traditional methods (McKinsey & Company,,2023).
While these AI tools are used either before the design generation stage to analyze consumer data and extract consumer needs or after the design generation stage to predict consumer demand for new designs, most of the current AI applications do not systemically incorporate large-scale consumer preference information into the design generation process. During the design generation stage, consumer preferences are integrated in a limited and ad hoc manner. For example, generative AI-based text-to-image software can be used to design new clothing of a certain style through iterative prompting (McKinsey & Company,, 2023). However, these prompts are not generated in a systematic way and may not capture consumers’ heterogeneous tastes. Even if new designs are generated automatically and at a low cost, how attractive they are to consumers remains unknown. To predict consumer demand for new designs, firms often need to employ A/B tests or “theme clinics” by asking consumers to evaluate different aesthetic designs. Theme clinics are costly; for example, firms usually spend more than \$100,000 conducting a single vehicle design test when designing cars (Burnap et al.,, 2023). In general, the task of predicting consumer demand is performed after generating new designs, which potentially wastes the resources spent developing and testing unpopular designs.
In contrast to the lack of incorporating consumer preferences into generative AI applications, rich consumer preference information exists in user-generated content (UGC) and remains underutilized by companies in practice. By nature, UGC data inherently contain rich consumer preference information as users voluntarily generate the associated content. For example, social media and UGC websites contain vast sets of photos taken by individual users in front of various backgrounds. The fact that a user chose to take a photo in front of a particular background means that the user liked the background. As such, UGC images can be used to understand how consumers evaluate new photo background designs. The backgrounds in UGC images can also serve as useful inputs for training generative models to generate new photo template designs. Despite the wide availability of UGC data, their usage in marketing is mostly descriptive and focuses on extracting consumer preference information or predicting consumer demand. Prescriptive UGC usage (i.e., designing products for companies) is scant both in marketing academia and in practice. No clear guidance has been provided on how to use unstructured UGC data to directly make unstructured firm decisions such as product design.
Another challenge lies in the cost of finding a scalable solution for incorporating consumer preferences and leveraging UGC data. The development of a generative AI algorithm may be costly and unaffordable for many companies. Training an algorithm also requires a large training dataset, which may not be available for small companies with limited internal data or start-up companies facing a “cold-start” problem.
Our study fills these gaps and addresses three key questions. (1) How can consumer preference information be incorporated into the product design process, and how can generative AI be guided to generate new designs that are appealing to consumers? (2) How can firms leverage the rich design and preference information contained in UGC during the design generation process, especially for firms facing “cold-start” problems? (3) How can the generation process be automated in a systematic, scalable, and cost-efficient way?
We propose a semi-supervised deep generative framework that creates consumer-preferred product designs in an automatic and cost-effective manner. The framework consists of a predictor model that learns consumer preferences and a generative model that uses the outputs of the prediction model to guide the generation process. This study has several highlights. First, different from previous studies in which consumer preference information was only used for training a predictor model, we incorporate consumer preference information into both our predictor and generator models. In particular, the predicted popularity for each product serves as a label and an additional input to the generator, which allows the generator to create products that are more likely to be preferred by consumers. Second, we leverage UGC acquired from external data sources. The various product designs contained in this UGC enrich the generative model and allow it to incorporate new product features that are not included in the internal data of firms. The rich consumer preference information embedded in the UGC improves the supervision of the generative model and allows it to generate appealing designs that go beyond the scope of the internal preferences and designs.
We present different versions of the generative model to illustrate the advantages of our proposed framework. We start with a simple GAN that does not incorporate any consumer preferences. We then gradually add consumer preferences to the generative process, first adding only internal preferences and then adding both internal and external preferences.
Specifically, the baseline model uses a deep convolutional GAN (DCGAN). Internal designs and external designs are used as inputs to generate new designs without incorporating consumer preferences. This approach extends the current industry practice by incorporating new designs from UGC. The second model uses the continuous conditional GAN (CcGAN) introduced by Ding et al., (2020). Compared to traditional GANs, the CcGAN can directionally generate new images of certain types. In particular, the CcGAN takes image labels as additional generative model inputs, which enables it to generate new images with a specific label. For example, a traditional GAN can use a set of dog images to produce a new set of dog images, while the CcGAN can instruct its model to create images of a specific type of dog (e.g., a golden retriever). The CcGAN is well suited for our task because it allows us to incorporate consumer preferences and generate new product designs that are more likely to be chosen by consumers. We first train a predictor model to learn consumer preferences from consumers’ historical choice data regarding existing designs. We then use the predicted choice probabilities acquired from the predictor model as additional input labels to guide the training process of the CcGAN. Finally, we instruct the trained CcGAN to generate new product designs that have higher predicted choice probabilities. We examine two versions of the CcGAN: an enhanced model and an advanced model. The enhanced model only uses internal consumer preference information to train the predictor model. The advanced model uses both internal and external consumer preference information to train the predictor model, and this variant represents the final version of our proposed framework.
We apply our proposed framework to a real business setting: a self-aided photography chain in China. The company operates self-aided photo booths in large shopping malls in major cities. The photo booths allow consumers to choose photo templates, take photos with the chosen templates, and print their photos. These photo templates are the main product offerings of the company, so it is crucial for the firm to design appealing photo templates to attract consumers. However, as a start-up company, the firm has limited resources and does not have a systematic way to design new templates. The firm also faces a “cold-start” problem with limited internal data. Our proposed framework can help the firm generate new consumer-preferred photo template designs in an automatic and cost-effective manner.
The internal dataset contains 2195 consumers who had taken at least one photo between October 2017 and August 2018. For each consumer, we obtain her full order history, including her selected photo templates and the photos taken with these chosen templates. We further collect external images from UGC websites. These external images are of similar types to those of the internal images, with humans taking photos in front of various backgrounds. The external dataset serves two purposes. First, it provides training inputs for the constructed generative models, allowing them to incorporate new product features derived from external product designs. It addresses the issue that the internal data contain limited style variations. Second, the rich history of external preferences inherent in the external data allows the predictor model to recognize popular product features from the external sources and guide the generative model to produce “out-of-the-box” appealing designs that do not exist in the internal data.
Comparing the performances of the generated templates produced by different models, we find that the advanced model with both internal and external consumer preferences performs best. The new designs generated by the advanced model are predicted to be more likely to be preferred by consumers by both the predictor model and outside metrics. Our results suggest that merely incorporating internal consumer preferences is not sufficient; it is important to further incorporate external consumer preferences, as external preferences provide additional information on how consumers evaluate external designs, thereby guiding the generative model to incorporate the popular features of the external designs that do not exist in the internal designs. Overall, the results highlight the importance of incorporating consumer preferences into the generation process.
Managerially, our work provides an automated product design solution for firms, which is especially valuable for smaller start-up firms that face cold-start problems and budget constraints. Firms have internal data on their existing product designs and the choices of consumers among the existing products. Our framework enables firms to further incorporate product design ideas and consumer tastes from external websites in a systematic, scalable, and cost-effective way. The proposed design process can also dynamically evolve as more internal and external data are incorporated as inputs, which enables the firms to rapidly respond to consumer demand shifts.
Our proposed model enhances the traditional product design process in three ways. First, it offers a superior alternative to the prevailing industry practice that relies solely on pure GAN models. The traditional methods do not consider consumer tastes during the design generation process; thus, generating a large set of new designs and screening unpopular designs on an ex post facto basis are needed. Our method generates new designs that are preferred by consumers ex ante, enabling companies to spend less time and resources on new designs that are later shown to be less appealing. Second, our proposed model provides an effective solution for product design ideation. The traditional product design ideation process relies on professionally trained human designers, who have limited capacity and may not systematically learn the dynamically evolving requirements of a large set of consumers. Our method leverages design ideas from UGC that reflect the real-time preferences of a large user base. Third, our proposed model offers a more cost-effective solution for obtaining consumer evaluations of new product designs. In contrast to traditional methods such as conducting costly consumer surveys, we leverage UGC as a source of real-time, self-revealed consumer preferences that are relevant and dynamically updating. Our method is applicable to other product design settings (e.g., video meeting backgrounds) where consumer preference information widely exists and external resources (e.g., UGC) are easily accessible.
2 Conceptual Model for Generating New Designs
2.1 Overview of the Conceptual Model
We propose a semi-supervised deep generative framework for generating consumer-preferred new product designs in a systematic and cost-efficient way. Our proposed model has three unique advantages. First, it allows firms to leverage product designs from external sources such as UGC on social media and image websites. This is especially beneficial to start-up firms with limited internal product designs. Second, it incorporates consumer preference information from the internal data of a firm during the design generation process so that the generated new designs are preferred by consumers. Third, it incorporates consumer preference from external sources, which allows firms to leverage the rich and widely available consumer preference information contained in UGC.
To illustrate these three advantages, we introduce three versions of the generative model. The first is a baseline model that uses product designs from both internal and external sources to generate new designs. This approach illustrates the first advantage of leveraging external product designs from UGC. The second variant builds on the baseline model and further incorporates consumer preference information acquired from the internal data of a firm. Compared with the first model, this version illustrates the second advantage of allowing internal consumer preferences to guide the generative process and directionally generate more appealing designs. The third variant is the full version that incorporates product design and consumer preference information from both internal and external sources. This method demonstrates all three advantages of our proposed framework.
In this subsection, we describe the intuitions and workflows of the three models. We discuss the implementation details of the models in the following subsections.
Figure 1: Proposed Machine Learning Model for Generating New Designs
Baseline model: DCGAN without consumer preferences
The baseline model uses existing product designs from both internal and external sources, as shown in Figure 1. The internal datasets of firms contain existing product designs that can be used as generative model inputs. However, these internal designs are usually limited in both quantity and style variety, especially for start-up firms that face cold-start problems. To overcome this difficulty, we gather existing product designs from external sources such as UGC on social media and image websites. UGC contains a rich set of images and image features. Utilizing UGC as inputs not only increases the size of the training dataset but also introduces new image features to the generative process; both of these factors enhance the performance of the generative model.
Studies have explored the use of UGC in marketing. Most of these studies were descriptive and focused on analyzing UGC to extract information and gain insights for firm operations. Some of them have analyzed textual content. For example, Netzer et al., (2012) extracted brand associations from UGC on forums. Culotta and Cutler, (2016) extracted the social connections of brands on Twitter. Other works analyzed visual content (Zhang and Luo,, 2023; Zhang et al.,, 2017; Pavlov and Mizik,, 2019) . For instance, Liu et al., (2020) analyzed consumer-generated images on social media to study the brand perceptions of consumers. We contribute to the literature by going one step further, from descriptive to prescriptive work. Our generative model not only extracts information from unstructured UGC but also uses this information to make unstructured product design decisions for firms.
Following industry standards, we use GANs (Goodfellow et al.,, 2014) to generate new product designs in the baseline model. We choose GAN models over other generative models such as fully visible belief networks (FVBNs) (Frey,, 1998) and variational autoencoders (VAEs) (Kingma and Welling,, 2013; Rezende et al.,, 2014). The reason for this is that GANs were designed to avoid the many disadvantages associated with other generative models and are regarded as producers of higher-quality samples than those of other methods. As we focus on image synthesis-based tasks, we use the DCGAN (Radford et al.,, 2015). The DCGAN is considered one of the most important steps forward in terms of designing and training stable GANs and has become the de facto standard given its high-level performance in image tasks.
Enhanced model: CcGAN with internal consumer preferences
Although the baseline model can generate new product designs in an automated, scalable, and efficient manner, it does not incorporate consumer preference information. Therefore, the newly generated designs may not fit consumer needs, and consumers may choose not to buy these new products. This is a general limitation encountered in most GAN applications in practice: GANs are used in an engineering sense without incorporating business insights and consumer needs.
We propose an enhanced generative model that incorporates consumer preference information into the generative process in an automated and large-scale manner. To our knowledge, this is one of the first studies to incorporate consumer preferences into a generative model. Most of the previous studies used consumer preference information only in prediction tasks. For example, Burnap et al., (2023) used consumer preference information to train a predictor model and applied the predictor model to predict how consumers perceived new product designs; the new product designs, however, were generated without incorporating consumer preference information. We differ in that we incorporate consumer preference information jointly in the predictor model and generator model. In particular, the predictor model captures consumer preferences and is used to guide the process of designing new products in the generator model. Therefore, the generator model can generate new products that are more likely to be preferred by consumers in a targeted manner.
The internal consumer preference information that we incorporate into the model comes from consumers’ historical product choices in the internal data of firms. Our empirical application is employed as an example. The internal data of the firm contains the photo templates that a particular consumer chose and the corresponding photos taken by the consumer with the chosen templates. Each photo taken contains both the consumer’s appearance and the chosen photo template. These consumer choices represent the types of product designs that a particular consumer prefers.
We first train a predictor model to learn the internal preferences of consumers, which we incorporate into the generative process. The predictor model utilizes the historical photos taken by internal consumers as inputs. It learns a mapping between a consumer’s appearance (i.e., her facial traits) and her preferred product designs, assuming that consumers with similar appearances will prefer similar product designs. To address the high dimensionality of the target problem, we preprocess the observed consumer faces and product designs using embedding models, obtaining low-dimensional vector representations for both consumer appearances and product designs. Using these embeddings as inputs, we train a deep prediction model to capture high-level abstractions in the prediction task. We call this predictor model with internal preferences “Prediction Model 1”, as illustrated in Figure 1. The trained predictor model can be used to predict the popularity of any given product design. Specifically, for each pair consisting of a product design and an internal consumer, the trained predictor calculates the probability that this consumer will choose the corresponding product. By averaging over the selection probabilities among all internal consumers, we can obtain a popularity measure of the particular product.
The trained predictor model is further integrated into the generator model in a semi-supervised deep learning framework. We employ a conditional GAN model (Mirza and Osindero,, 2014), which extends the traditional GAN by enabling the conditional generation of outputs. Intuitively, the traditional GAN can use a set of dog images to produce a new set of dog images, but one cannot instruct the model to create images of a specific type of dog. The conditional GAN allows one to condition the network on additional information such as class labels. During training, the conditional GAN takes images with labels (bulldogs, golden retrievers, etc.) as inputs and learns the differences between the images with different labels. In this way, it gains the ability to generate new images with a specific label (a specific type of dog). The conditional GAN is well suited for our task because it allows us to incorporate consumer preferences and generate new product designs that are more likely to be chosen by consumers. In our context, we use the popularity measures derived from the predictor model as additional inputs (i.e., class labels of product designs). The predicted popularity levels guide the training procedure of the conditional GAN so that internal consumer preferences are incorporated into the generative model. Once the generative model is trained, we can instruct the model to generate new product designs with higher popularity levels. In this way, the conditional GAN can generate new designs that are more appealing to consumers.
A key contribution of our proposed framework is its introduction of a conditional GAN to marketing applications. One challenge is that conditional GAN models typically use categorical labels as additional inputs, while the predicted popularity measures in our context are continuous class labels. To address this issue, we employ the CcGAN that was introduced by Ding et al., (2020). The CcGAN is an extension of the conditional GAN model, and it was designed to generate images based on continuous, scalar conditions. By redefining the conventional conditional GAN approach to make it suitable for continuous scenarios and introducing a novel method that can integrate continuous labels into the generation process, the CcGAN outperforms the conditional GAN in terms of effectively handling continuous labels.
Advanced model: CcGAN with internal and external consumer preferences
One limitation of the enhanced model is that the predictor model is trained only on internal consumer choices and lacks exposure to diverse design styles that are not present in the internal data. Consequently, the predictor model may struggle to understand how consumers evaluate new design styles from the external product designs. Even if the generative model is trained on both internal and external product designs and is able to produce “out-of-the-box” styles (i.e., styles that are not present in the internal designs), the predictor model may not recognize their attractiveness, which limits the performance of the CcGAN.
Another limitation of the enhanced model is that it does not fully utilize the information contained in external UGC. By nature, UGC data inherently contain rich consumer preference information, as users voluntarily generate the content. In our context, i.e., photo templates, the UGC on social media and image websites usually contains a vast set of photos taken by individual users in front of various backgrounds. This type of image represents not only a rich trove of product design data (i.e., the backgrounds in the case of photo templates) but also a valuable source of consumer preference information. The fact that an individual user chose to take photos in front of a particular background means that the user likes the background. As such, the external images can also be used to train the predictor model to understand how consumers evaluate new features in external product designs.
We propose an advanced model that integrates both internal and external consumer preferences when generating new product designs. While the previous studies used external images as extra inputs for training their generative models (Burnap et al.,, 2023), we further exploit the consumer preferences revealed in the external data to train both the predictive and generative models. Therefore, we can make the most of the available UGC since it contains rich consumer preference information. Such an approach also saves firms the significant time and money they would have spent on conducting ex post consumer product tests. As discussed in Burnap et al., (2023), firms often employ A/B tests, or “theme clinics”, to ask consumers to evaluate different aesthetic designs to improve their product designs. These theme clinics are costly, as firms usually spend more than \$100,000 to conduct a single vehicle design test. Therefore, utilizing consumer preferences derived from external UGC is especially useful for start-up firms with limited resources. By directly using the existing consumer preference information contained in UGC, our proposed model significantly reduces the cost of gathering consumer preference information.
The advanced model follows a workflow similar to that of the enhanced model except for the fact that the predictor model is trained on both internal and external consumer preferences. Following the same logic as that used for the internal data, the predictor model learns the mappings between the users’ appearances and the backgrounds in the UGC images. Upon pooling the mappings between the consumer appearances and product designs in both the internal and external datasets, the trained prediction model captures a larger set of consumers and their preferences over a wider variety of product designs. The predictor model is labeled Prediction Model 2 in Figure 1. Compared with Prediction Model 1, this version can recognize popular out-of-sample features from external designs.
The remainder of the workflow is the same as that of the enhanced model. We use the trained predictor model to compute average choice probabilities for each product. The average choice probability of a product serves as its popularity measure and is then used as a continuous label in the CcGAN, guiding the model to generate popular images. Importantly, the new predictor model overcomes the limitations of the previous predictor model, which only recognizes popular product features within internal consumer preferences. The new popularity measure represents a wider range of consumer preferences over a wider range of product designs. Consequently, the CcGAN with both internal and external consumer preferences is able to generate new product designs that better incorporate the new features acquired from the external data.
As shown in Figure 1, the external dataset serves two purposes. First, the product features obtained from the external data enhance the training processes of GAN models. This addresses the issue regarding the limited style variations contained in the internal data by introducing additional style diversity to train the GAN models to generate improved designs. Second, the rich trove of external preferences inherent in the external data guides the GAN models to generate more appealing designs that go beyond the scope of the internal preferences and designs.
Summary of the models
Comparing the three models illustrates the benefits of incorporating external designs and consumer preference information into the product design generation process. The full version of our proposed model (the CcGAN with internal and external consumer preferences) is a semi-supervised deep generative framework that creates consumer-preferred product designs in an automatic and cost-effective manner. It has several highlights. First, different from previous studies in which consumer preference information was only used to train predictor models, we incorporate consumer preference information into both the predictor model and the generator model. In particular, the popularity predicted for each product serves as a label and an additional input for the generator, which allows the generator to create products that are more likely to be preferred by the target consumers. Second, we exploit the UGC derived from external data sources. The various product designs contained in UGC enrich the generative model and allow it to incorporate new product features that are not included in the internal data. The rich consumer preference information embedded in the UGC improves the supervision of the generative model and allows it to recognize popular new product features that are not presented to the internal consumers.
We contribute to the literature concerning deep learning models in marketing. Most studies in this area use unstructured data to generate structured firm insights. For instance, Liu et al., (2020) collected unstructured image data from social media and used a deep convolutional neural network (CNN) to extract and predict consumer perceptions of brand attributes. Our work differs in that we employ unstructured data to directly generate unstructured firm decisions (i.e., product designs as images) using state-of-the-art deep learning models. Our proposed framework goes beyond the descriptive task and provides prescriptive solutions for the product design decisions of firms.
We also contribute to the nascent literature on deep generative models in marketing. A few works have used deep generative models to create new product designs. For example, Dew et al., 2022 used multiview representation learning to design brand logos. Burnap et al., 2023 proposed a model to augment the commonly used aesthetic design process by predicting aesthetic scores and automatically generating innovative and appealing product designs; the predictor model and the generator models were separately trained. Our work differs in that we incorporate consumer preference information into the generation process so that the predictor model guides the generative model. We also introduce the CcGAN model to the marketing field, allowing us to systematically incorporate consumer preferences in an automated and large-scale manner. In addition, while previous works generated one aspect of product designs such as logos (Dew et al.,, 2022) or the aesthetic appearances of cars (Burnap et al.,, 2023), we directly design the entire product (i.e., the photo templates are the products of the firm). This setting allows us to abstract away from other design factors that might influence consumer choices.
2.2 Prediction Model
This subsection describes the detailed procedure for training the prediction model, as shown in Equation 2.1. The primary inputs of the prediction model consist of consumer appearances (i.e., consumer faces in our empirical context) and product designs (i.e., photo templates or backgrounds in our empirical context). Given the high dimensionality of the input data, we preprocess the given images by embedding them into a lower-dimensional vector space. For consumer faces, we follow the deep neural network implemented by the OpenFace project (Amos et al.,, 2016). As illustrated in Figure A.1 in the appendix, this architecture was trained for facial recognition. It first detects and crops faces from a given image. Then, it produces a 128-dimensional intermediate layer through a deep neutral network, which represents a low-dimensional embedding of the face image. For the product design images, we adopt the Caffe BVLC reference model (Jia et al.,, 2014) to obtain image embeddings. This model uses the CNN designed by Alex Krizhevsky (AlexNet) with some modifications, as shown in Figure A.2 in the appendix. It provides pretrained general networks that can be used to obtain low-dimensional embeddings of image characteristics. Using the pretrained weights of the Caffe BVLC reference model, we can directly go to fully connected (fc) layer 8 to obtain a 1000-dimensional intermediate layer that captures the visual characteristics of the input product design image.
[…]
Let the vector $\overrightarrow{X}_{i}$ denote the 128-dimensional embedding of consumer $i$’s face. Let the vector $\overrightarrow{V}_{j}$ denote the 1000-dimensional embedding of product design $j$. Let the vector $\overrightarrow{Z}_{j}$ denote any other low-dimensional product characteristics that affect consumer’s product choices. We concatenate these three vectors to form the final input vector for training the predictor model. Let $Y_{ij}$ denote the outcome variable, or the choice outcome for consumer-product pair $(i,j)$. $Y_{ij}=1$ if consumer $i$ chooses product $j$, and $Y_{ij}=0$ otherwise. The predictor model is trained to learn the nonlinear relationship between the input and output variables:
$Y_{ij}=f(\overrightarrow{X}_{i},\overrightarrow{V}_{j},\overrightarrow{Z}_{j}|\theta,\varepsilon),$
where $\theta$ is the model parameter to be estimated. Since the prediction task is a classification problem with relatively high-dimensional inputs, we adopt a random forest to estimate the model parameter $\theta$. After obtaining the trained predictor, we can use it to calculate the probability that consumer $i$ will choose product $j$:
$p_{ij}(\widehat{\theta})=\mbox{Pr}(Y_{ij}=1|\overrightarrow{X}_{i},\overrightarrow{V}_{j},\overrightarrow{Z}_{j};\widehat{\theta})$
By averaging over the choice probabilities of all consumers in the sample, we can obtain an aggregate popularity measure for product $j$:
$P_{j}(\widehat{\theta})=\frac{1}{N}\sum_{i\in\mathcal{I}}p_{ij}(\widehat{\theta}),$ (2.2)
where $\mathcal{I}$ denotes the total consumer set and $N$ is the size of the consumer base. The aggregate popularity measure is later incorporated into the constructed generative models to guide the process of generating new product designs.
As discussed in Section 2.1, we train two prediction models: one with internal preferences only and another with both internal and external preferences. We train these two prediction models via similar training processes. The difference is that Prediction Model 1 is trained with internal consumer appearances and product designs only, while Prediction Model 2 is trained using both internal and external consumer appearances and product designs. Note that external consumers are used in the training stage of Prediction Model 2 but not in the prediction stage; when using the trained models to calculate the aggregate popularity measures, we average over the choice probabilities of the internal consumers only for both Prediction Models 1 and 2. This is because we are interested in predicting whether the internal consumers of firms prefer the new template designs. The external consumer preference information still enhances the performance of Prediction Model 2 by training the model to recognize popular design features in external designs. The underlying assumption is that external consumers with similar appearances to those of the internal consumers share similar preferences. Therefore, learning the preferences of the external consumers helps us infer the preferences of the internal consumers for the new design features acquired from the external designs.
2.3 Generative Models
This subsection describes the detailed procedure for training the generative models.
When training the baseline model without consumer preferences, we adopt the standard and well-established model architecture of the DCGAN method (Radford et al.,,2015) to generate new images. This model takes the internal and external product images as real images input and identifies a set of parameters that enables the samples generated from random noises to closely resemble the training data. As the training process is unsupervised and lacks consumer preference information, the generated images are not necessarily preferred by consumers.
When training the enhanced model with internal consumer preferences, we estimate a conditional GAN model and let the internal consumer preferences guide the generator to produce appealing designs that are tailored to the targeted consumers. Specifically, we use the CcGAN developed in Ding et al., (2020) to handle continuous labels.
Let $M^{r}$ and $M^{g}$ denote the numbers of real and generated Images, respectively. Let $\boldsymbol{x}\in\mathcal{X}$ denote an image with a size of $d\times d$. Let $y\in\mathcal{Y}$ denote the corresponding label. We pool together the existing internal and external product designs to form a group of real images as input images, yielding a sample size of $M^{r}$. For each image $j$ in the input group, we use Prediction Model 1 to calculate its aggregate popularity measure:
$y_{j}=P_{j}(\widehat{\theta}_{1}),$ (2.3)
where $\widehat{\theta}_{1}$ is the estimated parameter of Prediction Model 1 and $P_{j}(\widehat{\theta}_{1})$ is defined in Equation 2.2. This popularity measure serves as a continuous class label and is fed into both the generator and the discriminator in the CcGAN to guide the generator when creating new designs. Specifically, following Ding et al., (2020), we define the hard vicinal discriminator loss (HVDL) as
^$\begin{split}\widehat{\mathcal{L}}(D)=&-\frac{C_{1}}{M^{r}}\sum_{j=1}^{M^{r}}\sum_{i=1}^{M^{r}}\mathbb{E}_{\epsilon^{r}\sim N(0,\sigma^{2})}\Big{[}\frac{\mathbbm{1}_{\{|y_{j}^{r}+\epsilon^{r}-y_{i}^{r}|\leq\kappa\}}}{N_{y_{j}^{r}+\epsilon^{r},\kappa}^{r}}\mbox{log}(D(\boldsymbol{x}_{i}^{r},y_{j}^{r}+\epsilon^{r}))\Big{]}\\ &-\frac{C_{2}}{M^{g}}\sum_{j=1}^{M^{g}}\sum_{i=1}^{M^{g}}\mathbb{E}_{\epsilon^{g}\sim N(0,\sigma^{2})}\Big{[}\frac{\mathbbm{1}_{\{|y_{j}^{g}+\epsilon^{g}-y_{i}^{g}|\leq\kappa\}}}{N_{y_{j}^{g}+\epsilon^{g},\kappa}^{r}}\mbox{log}(1-D(\boldsymbol{x}_{i}^{g},y_{j}^{g}+\epsilon^{g}))\Big{]},\end{split}$ (2.4)
where $C_{1}$ and $C_{2}$ are constants, $\boldsymbol{x}_{i}^{r}$ and $\boldsymbol{x}_{i}^{g}$ stand for real and generated images, respectively, $y_{i}^{r}$ and $y_{i}^{g}$ are the popularity labels of $\boldsymbol{x}_{i}^{r}$ and $\boldsymbol{x}_{i}^{g}$, and $\epsilon^{r}=y-y_{j}^{r}$ and $\epsilon^{g}=y-y_{j}^{g}$. $N_{y_{j}^{r}+\epsilon^{r},k}^{r}$ is the number of the $y_{i}^{r}$ that satisfies $|y_{j}^{r}+\epsilon^{r}-y_{i}^{r}|\leq\kappa$. The loss function for the generator is defined as
^$\widehat{\mathcal{L}}(G)=-\frac{1}{M^{g}}\sum_{i=1}^{M^{g}}\mathbb{E}_{\epsilon^{g}\sim N(0,\sigma^{2})}\mbox{log}(D(G(\boldsymbol{z}_{i},y_{i}^{g}+\epsilon^{g}),y_{i}^{g}+\epsilon^{g})),$ (2.5)
where $\boldsymbol{z}_{i}$ is random noise that follows a normal distribution $N(0,\boldsymbol{I})$. As demonstrated in Ding et al., (2020), employing the HVDL allows us to train the discriminator using the images near $y$ rather then solely relying on images labeled $y$. This approach addresses the scarcity of real images for each label value. Additionally, we adopt the innovative input method in Ding et al., (2020) to incorporate labels into the model. Specifically, in the generator, we add the labels to the feature map of the first linear layer in an elementwise manner. In the discriminator, we first project the labels to a latent space through an additional linear layer. We then integrate these embedded labels into the discriminator via label projection.
In summary, combining Equations 2.3, 2.4, and 2.5, we integrate consumer preferences into the generation process by directly incorporating the popularity measure $P_{j}(\widehat{\theta})$ from the predictor model into the loss functions of both the discriminator and the generator of the CcGAN. This setup allows us to instruct the generator to create new images with a specific popularity level.
The training process of advanced model with both internal and external consumer preferences follows a procedure similar to that used for training the enhanced model. The only difference is that we use the aggregate popularity measures from Prediction Model 2, $P_{j}(\widehat{\theta}_{2})$, as additional input labels to be fed into the generator and the discriminator. As Prediction Model 2 is trained on both internal and external consumer choices, the new CcGAN is able to incorporate new design features from the external sources that are appealing to the internal consumers.
3 Institutional Setting and Data
3.1 Institutional Background
As an empirical application, we apply our proposed framework to a real business setting: a large self-executed photography chain in China. The company operates self-executed photo booths in approximately twenty shopping malls in major cities. The photo booths allow consumers to choose photo templates, take photos with the chosen templates, and print their photos. The photo booths are self-operated, so consumers follow the instructions on the photo machines placed in the booth. Figure 2 shows the interface of such a photo machine. Four photo categories can be selected: main (single), friends, family, and lovers. We focus on the “main” category in our analysis, which is designed for one consumer to take photos.
The decision process of the consumers is as follows. Upon entering the photo booth, the consumers are asked to choose photo themes. Several pages of themes are available for the consumers to choose from. The themes range from general topics such as “modern” and “classic” to very specific topics such as “fruits” and “flowers”. The company constantly introduces new themes and ranks these themes by recency, with the more recent themes placed on the first pages on the screen. After clicking on a specific theme, the consumers are presented with all templates under this theme in a pop-out window, and they choose the template(s) with which they will take photos. Each template is typically a combination of a background and a foreground without human subjects involved. For example, Figure 3 shows four photo templates, two from the “traditional Chinese” theme and two from the “graffiti” theme. The templates within a specific theme share a similar style, while templates across themes can differ substantially from each other. After choosing the templates, consumers pay for a time window (10, 30, or 60 minutes) and use the time to take photos with the chosen templates. The company charges consumers based on the time they use to take photos, not the number of templates chosen. Consumers can choose an unlimited number of templates and take photos as long as they finish within the paid time window.
Figure 2: Interface of the Self-Operated Photography Booth
Figure 3: Examples of Photo Templates
Refer to caption Refer to caption
(a) Graffiti Theme (b) Traditional Chinese Theme
As it is in its early development stage, the company does not have a systematic and cost-effective way to design new themes or new templates due to internal resource limitations. The company hires outside designers to create new designs without incorporating consumers’ aesthetic preferences. The process of ideating the new themes and designs is not systematic, either, except that holiday-related themes are introduced during holiday seasons. The company manages existing and new designs in a relatively unsophisticated and ad hoc manner: they introduce new designs at a constant pace and remove existing themes that turn out to be unpopular. Our proposed framework can help the company generate new designs in a systematic, cost-effective, and automated way. In particular, the generation process incorporates consumers’ heterogeneous preferences and leverages valuable designs and preference information derived from external sources.
3.2 Internal Data
The internal dataset obtained from the company contains two parts. The first part includes the original template images that the company offered to customers, without the presence of consumers inside the images. The company offered 83 photo themes in total during the observation period. Each theme contains 1 to 12 templates, yielding a total of 585 templates across all the themes. The second part of the internal dataset contains the complete order histories of 2195 consumers who had taken at least one photo between October 2017 and August 2018. For each consumer, we observe their selected themes and templates and the photos taken with the chosen templates. We also observe the ranking orders of the themes offered on the screen at the times they made their choices.
The internal data patterns suggest that the templates vary significantly in popularity, potentially due to different design features and heterogeneous consumer preferences. Figure 4 presents the number of consumers who choose each template in the graffiti theme and the traditional Chinese theme. The X-axis is the template index, which represents the order in which the templates belonging to each theme are displayed on the screen. As shown in the figure, the templates have very different popularity levels both across themes and within the same theme. For graffiti, the number of adoptions for the most-popular template is four times that of the least-popular template. For the traditional Chinese theme, the least-preferred template has zero adoptions, whereas the most-preferred template has more than 60 adoptions. The variation in template popularity, especially the presence of low-popularity templates, suggests that the company needs to improve the attractiveness of its template designs. As another piece of evidence, the internal dataset shows that only 2195 out of the 6038 registered consumers ended up taking photos in the booth after browsing the photo template options. Among the 2195 consumers, fewer than 100 consumers made multiple store visits, and only 15% of the consumers had repeated time purchases. These numbers suggest that the company can benefit from incorporating consumer preferences to generate better designs that are appealing to consumers.
Figure 4: Number of Adopters of Each Template Under Different Themes
Refer to caption Refer to caption
(a) Graffiti Theme (b) Traditional Chinese Theme
3.3 External Data
To leverage the existing designs and consumer preference information derived from external sources, we further collect an external dataset from two main sources: social media and online image websites. Social media reflects various aspects of consumer tastes, whereas image websites contain a variety of photos with different themes. We collect specific types of photos from these websites, i.e., photos taken by individual users in front of specific types of backgrounds, which are similar to the photo-taking setting in our case. The backgrounds in these external photos can serve as additional design templates for training the generative model. The mappings between the individuals and the backgrounds in these photos can serve as additional consumer preference information to train the prediction model.
After gathering the external images with human objects and backgrounds, we preprocess the images before they are used as inputs for the proposed model. Specifically, training the model requires a set of pure templates without human objects and a set of consumer face images, while the raw external images we collect are combinations of template backgrounds and consumer faces, so we need to separate the two components. For each external image, we remove the human object in it and inpaint the remaining part of the image to obtain a complete background image, which is similar to the pure template images contained in the internal dataset. Additionally, we save the cropped human faces as new images, which are similar to the consumer face images in the internal dataset. Figure 5 presents examples of the internal templates, raw external photos, inpainted template images, and cropped consumer face images, all of which are from the graffiti theme. As shown in the graphs, the collected external images are similar in style to the internal images and have sufficient variations. The inpainting process works well and can generate high-quality templates and consumer face images.
Figure 5: Examples of Graffiti Templates: Internal vs. External
Refer to caption Refer to caption
(a) Internal images (b) External images (raw, inpainted, cropped)
4 Applying the Model to Practice
We apply the proposed model to generate new product designs for the unmanned photo gallery firm. This section describes the implementation details, and the next section presents the results.
4.1 Training the Prediction Model
As discussed in Section 2, the feature embeddings of consumer faces and background templates serve as the inputs for the prediction model. Regarding the consumer face features, most consumers took more than one photo, so multiple face images are available for each consumer. We first obtain the embedding of each face image. We then average across the embeddings of the same consumer to obtain a 128-dimensional facial vector $\overrightarrow{X}_{i}$ for that consumer. With respect to the template features, in addition to the 1000-dimensional embeddings of the visual characteristics $\overrightarrow{V}_{j}$, we also include the display ranking of each template on the screen $\overrightarrow{Z}_{j}$, as its display ranking affects its probability of being chosen. Concatenating $\overrightarrow{X}_{i}$, $\overrightarrow{V}_{j}$, and $\overrightarrow{Z}_{j}$ yields a 1129-dimensional feature input for training the prediction model.
The output variable of the prediction model is constructed from the consumers’ revealed preferences, namely, their template choices. In our setting, consumers first choose a theme among all themes; only one template from each theme is displayed as the cover picture for each theme at this stage. Once consumers click on the chosen theme, they can browse all templates in the chosen theme and choose some of the templates to take photos. We assume that consumers make binary choices for each template they browse and can choose multiple templates since consumers can select an unlimited number of themes and templates within the paid time window. The chosen templates belonging to the chosen theme naturally serve as the positive samples. The unchosen templates belonging to the chosen theme serve as the negative samples because consumers see them but decide not to take photos with them. For the unchosen themes, the cover-page templates count toward the negative samples because consumers see them upon choosing the themes and do not choose them. However, the other templates belonging to the unchosen themes do not count as negative samples and are not included in the analysis because consumers do not have a chance to see them. Overall, the output variable $Y_{ij}$ for consumer $i$ and template $j$ is constructed as follows:
$Y_{ij}=\left\{\begin{array}[]{ll}1&\text{if template }j\text{ is chosen by consumer }i\\ 0&\text{if template }j\text{ is not chosen but belongs to a chosen theme}\\ 0&\text{if template }j\text{ is the cover-page template of an unchosen theme }\end{array}\right..$ (4.1)
Given the definitions of the input and output variables, we estimate Prediction Model 1 using internal consumer preferences only. We focus on the photo samples from the category “Main (single)”, which contains 931 consumers and 33,804 consumer-template pairs. Each consumer-template pair corresponds to one consumer choice. Among these 33,804 choices, 3,629 are positive samples (i.e., consumers choose to take photos with the templates), and 30,175 are negative samples. We retain $40\%$ of the observations as the test set and use the remaining data for training the model. The training and test sets include 20,282 and 13,522 observations, respectively. As positive samples only account for 10.7% of the total sample size, an unbalanced sample problem is encountered. To alleviate this problem, we upsample the minority class to 18,000 to better match the majority class. We concatenate the upsampled minority class and the majority class to form the final training set and train the model using a random forest. The number of trees in the forest is set to 100. The class weights are set to be balanced.
We estimate Prediction Model 2 using both internal and external consumer preferences. Specifically, external UGC data contain photos taken by individual users in front of various backgrounds. Each photo represents a user’s chosen background and thus serves as a positive sample for the prediction model. There are 1375 external photos or choice observations. Combining these with the 33,804 internal observations yields a sample size of 35,179 photos. We use 70% of the observations, or 24,625 photos, as the training set. Among them, 2,544 are positive samples, and 22,081 are negative samples. To address the unbalanced sample issue, we upsample the minority class to 20,000 to align with the majority class. We estimate Prediction Model 2 using a random forest again, maintaining the same hyperparameters as those employed in Prediction Model 1.
4.2 Training the Generative Models and Generating New Designs
Given the trained prediction model, we proceed to generate new designs using the three generative model versions and compare their performances. For illustrative purposes, we focus on generating new designs for the graffiti theme, as it is one of the most popular themes in the data. We present the results obtained for the traditional Chinese theme in Appendix B to show the generalizability of our method to different themes.
Baseline model: DCGAN without consumer preferences
To train the DCGAN model, we combine the existing templates acquired from the internal and external sources, resulting in 1388 training templates for the graffiti theme, 13 of which are from internal data. We exclude images with extreme height or width values. The final training set contains 1321 images. Since the size of our training data is relatively small and the performance of GANs heavily deteriorates in this case, we adopt the differential augmentation (DiffAugment) method proposed by Zhao et al., (2020). As the authors pointed out, the performance of a GAN depends heavily on the sample size of the input training data because the discriminator will be simply memorizing the exact training set given the small sample size. The DiffAugment method imposes various types of differentiable augmentation on both real and generated samples. It is able to generate high-fidelity images using only 100 images without pretraining, which sufficiently trains the model using a relatively small training set.
We produce output images with dimensions of $3\times 128\times 128$. The latent dimensionality is set to $100$, as is the standard value for the DCGAN. We use a batch size of $32$ for training the model, with the learning rate set to $2\times e^{-4}$, and the sizes of the feature maps in the generator and discriminator are both 64. To obtain stable results, we train the DCGAN for $5000$ epochs, implementing $275,000$ iterations in total.
Enhanced model: CcGAN with internal consumer preferences
To train the CcGAN with internal consumer preferences, we first use Prediction Model 1 to calculate the aggregate popularity levels for each of the 1388 existing templates. One caveat is that the prediction model requires the display rankings of the templates as inputs. We observe the rankings only for the internal templates but not for the external templates, as the external templates are from UGC websites and are not displayed in the system of the firm. We address this issue by simulating rankings for the external templates based on the observed empirical distribution of the rankings of the internal templates. The remainder of the procedure for calculating the popularity levels follows that discussed in Section 2.3.
Training the CcGAN requires a sufficient number of images within each class label. In our case, the popularity level takes continuous values, so the number of images with exactly the same popularity level is small. Therefore, we group the images with similar popularity levels into bins and let the images belonging to the same bin share the same popularity label. We focus on the images with popularity labels within the range of 0.4 to 0.65 since the majority of the images fall within this range. We limit the maximum number of images per label to 200 to produce a more balanced distribution. This results in a dataset containing 1317 images across 11 popularity labels. Figure A.3 in the appendix presents histograms for both the original popularity levels and the binned popularity levels. Finally, to facilitate comparisons across generative models, we normalize each label by dividing it by the maximum label value, resulting in 11 normalized labels ranging from 0.6154 to 1. These normalized labels are then used as continuous labels to guide the new design generation process.
We use the HVDL as the discriminator loss, where the specific loss functions are defined in Equations 2.4 and 2.5. Regarding the hyperparameters of the HVDL, we use the rule-of-thumb formulae to select them, i.e., $\sigma\approx 0.028$ and $\kappa\approx 0.192$. The CcGAN is trained for 20,000 iterations with a constant learning rate of $10^{-4}$ and a batch size of 64. The dimensions of the latent GAN are set to 256, and the discriminator updates twice in one iteration. To solve the small sample size problem, we adopt the DiffAugment method during the training process, similar to what we did in the DCGAN case. After training the CcGAN, we instruct the generator to produce new template designs with different popularity labels, where label=0.6154 indicates the least-preferred class, and label=1 indicates the most-preferred class.
Advanced model: CcGAN with internal and external consumer preferences
Following a procedure similar to that used for training the enhanced model, we use Prediction Model 2 to calculate popularity levels for the existing templates. We focus on images with popularity labels between 0.65 and 0.9 since the majority of the images fall into this range, and we cap the maximum number of images per label at 200 for balance purposes. This process results in 1283 images, which are then categorized into 10 labels with normalized label values between 0.7429 and 1. Figure A.4 in the Appendix illustrates the original and grouped popularity levels. The hyperparameters of the HVDL, $\sigma\approx 0.019$ and $\kappa\approx 0.143$, follow the rule-of-thumb formulae.
5 Results
In this section, we provide the results of our empirical application involving the unmanned photo gallery. Section 5.1 presents performance tests concerning the prediction models, including the prediction model with internal preferences only and the prediction model with both internal and external preferences. Section 5.2 shows the validity test conducted for the three generative models, including the DCGAN, the CcGAN model with internal preferences, and the CcGAN model with both internal and external preferences. Section 5.3 compares the performances of the three generative models.
5.1 Performance of the Prediction Models
In this section, we present performance tests concerning our prediction models. As the prediction models are trained on templates from all themes, the results are also obtained on all the themes.
Predicting actual consumer choices
We first show that our prediction model is able to predict observed consumer choices. Figure 6 presents examples of the predicted and actual choices of two consumers in the held-out sample. The first column shows the model-predicted consumer choices, which are defined as having predicted selection probabilities exceeding $0.5$. The second column shows the actual choices and the photos taken by the consumer. We can see that the actual choices are among the choices predicted by the model, but some predicted templates were not chosen by the consumers (i.e., false positives). [^2] The false positives in this illustrative test are not a major concern because they change with the threshold that we set. Here, we set a relative conservative threshold of $0.5$ by defining a model-predicted choice as having a predicted selection probability exceeding $0.5$. Setting a larger threshold will reduce the number of false positives. In the generative model, we use the continuous predicted choice probabilities as popularity labels, which are not affected by the threshold we set here. These examples provide direct evidence that our model can capture consumers’ heterogeneous and nonlinear preferences for the template designs fairly well.
Figure 6: Examples of Predicted and Actual Consumer Choices
Consumer Templates with predicted probabilities $>0.5$ Templates chosen by the consumer
Consumer 1 Refer to caption Refer to caption
Consumer 2 Refer to caption Refer to caption
Figure 7 presents the observed and predicted numbers of adoptions for three internal templates in the test set. We can see that the true and predicted numbers of adoptions are quite close to each other, suggesting that our prediction model is able to capture consumer preferences for different template designs.
Figure 7: Examples of the Predicted Numbers of Template Adoptions
Design popularity prediction
We then conduct a preliminary visual assessment to show that our prediction model is able to predict the popularity levels of different template designs. We identify the internal templates that are most (least) preferred by consumers based on their observed numbers of adoptions (row 1 of Figure 8). We then compare their styles with those of the external templates that our model predicts to have the highest (lowest) average choice probabilities (row 2 using Prediction Model 1 and row 3 using Prediction Model 2). We find that the external templates with high (low) predicted choice probabilities exhibit similar styles to the most (least) popular internal templates, suggesting that the predicted popularity levels are valid and accurate.
Figure 8: Comparing the Observed and Predicted Popularity Levels
Most-Preferred Templates Least-Preferred Templates
Observed Refer to caption Refer to caption
Prediction Model 1 Refer to caption Refer to caption
Prediction Model 2 Refer to caption Refer to caption
Prediction accuracy tests against the benchmark models
Next, we conduct prediction accuracy tests for Prediction Model 1 and Prediction Model 2 on the test set.
In the first row of Table 1, Prediction Model 1 has an overall accuracy rate of $0.793$. Given the unbalanced nature of our sample, another important performance measure is the balanced accuracy rate. The balanced accuracy rate of our model is $0.791$, which is almost as high as the overall accuracy rate. We further check the false-negative rate (FNR) and the false-positive rate (FPR) of our model and obtain an FNR of 0.212 and an FPR of $0.206$. These prediction test results show that our prediction model is able to use consumers’ facial traits to predict their template choices at a reasonably high accuracy level.
We also find that incorporating external preferences can help increase the overall prediction accuracy of the model. As shown in the second row of Table 1, the overall accuracy rate of Prediction Model 2 increases to 0.806, and its balanced accuracy rate is even higher, at 0.825. Both the false-negative rate and the false-positive rate are lower.
Finally, we compare our method with the logistic lasso model, which is a commonly used benchmark for binary outcome models. We choose logistic regression with lasso regularization for two main reasons. First, logistic regression is a prevalent linear and parsimonious model for classification tasks. Second, the lasso model is suitable for our context because it helps address the high dimensionality problem of our input variables (i.e., 1001+128+1) by automatically selecting relevant features and penalizing the inclusion of numerous trivial predictors.
Table 1 compares the results of the two models. Our proposed model outperforms the benchmark model with higher accuracy and balanced accuracy rates. This confirms the superiority of our proposed model in terms of predicting consumer template choices.
Table 1: Prediction Accuracy Achieved on the Test Set
Prediction Model 1 Prediction Model 2 Benchmark
Training Acc. Balanced FNR FPR Acc. Balanced FNR FPR Acc. Balanced
Input Acc. Acc. Acc.
Internal pref. 0.793 0.791 0.212 0.206 0.767 0.789
only
Internal and 0.806 0.825 0.148 0.202 0.78 0.815
external pref.
Note: This table presents the prediction accuracies, balanced accuracy rates, false-negative rates, and false-positive rates achieved on the test set using Prediction Model 1, Prediction Model 2, and the benchmark logistic lasso model. Prediction Model 1 is trained on the internal preferences only. Prediction Model 2 is trained on both the internal and external preferences. The benchmark model is trained on the internal preferences in the first row and both the internal and external preferences in the second row.
5.2 Validity of the Generative Models
After presenting the performance test for the prediction models, we now implement performance tests for the generative models. In this subsection, we demonstrate the face validity of the generative models by showing their ability to generate meaningful product designs with certain popularity levels. In the next subsection, we compare the performances of the generative models by comparing their abilities to generate product designs that are more likely to be preferred by consumers.
Face validity: Generating meaningful designs
We first show that our generative models are able to generate meaningful new designs. Intuitively, traditional GAN models can use a set of dog images to produce a new set of dog images. In our case, we need to show that our new model-generated designs are derived from the same theme as that of the input designs (i.e., graffiti). We start with a simple visual assessment. Figure 9(b) compares the real images and the new images generated by the trained DCGAN. We can see that the generated images are similar in style to the real images. At the same time, the generated images exhibit sufficient style variations, which guarantees that no modal collapse occurs during the training process.
Figure 9: Real vs. DCGAN-Generated Images
(a) Real
(b) DCGAN-Generated
In addition to this simple visual examination, we use a pretrained image classifier, ResNet18 (He et al.,, 2016), to conduct transfer learning to validate the meaningfulness of our DCGAN-generated images. Intuitively, if the generated images are indeed from the same theme as that of the real images (i.e., graffiti), ResNet18 should be able to classify them into the same category as that of the real images. First, we retrain the classifier to recognize graffiti and nongraffiti images using 100 real graffiti images (both internal and external, with $\text{label}=1$) and 70 real nongraffiti images (internal, with $\text{label}=0$) as inputs. Second, we test the validity of the retrained classifier on a test set consisting of real images. As shown in Table 2, the classifier has an accuracy rate of 97.2% when classifying the images into graffiti and nongraffiti categories. The false-negative rate and false-positive rate are both very low on the test set, further validating the ability of the classifier to predict the graffiti category. Finally, we use this retrained classifier to predict the labels for the set of generated graffiti images and present the hit rate in the last column of Table 2. Here, the hit rate is defined as the percentage of the generated images that are predicted to be in the graffiti category (i.e., $\text{label}=1$). We find that the hit rate is 99%, meaning that almost all the generated images are classified into the graffiti category by the classifier. This means that our generative model is able to produce meaningful designs with high fidelity.
Table 2: Image Classification Test Concerning the Generated Images
Test Set of Real Graffiti Images Generated Images
Classification Model Accuracy FNR FPR Hit Rate
ResNet18 Transfer Learning 0.972 0.024 0.034 0.99
Controllably generating designs with certain popularity levels
One advantage of our proposed framework is that it can be instructed to generate designs with a certain popularity level. This contrasts with traditional GAN models, which cannot be instructed to create images of a specific type. We start with the CcGAN model that incorporates internal consumer preferences. To demonstrate its ability to generate new designs with different popularity levels, we instruct it to generate new designs from the least-preferred class and the most-preferred class. The first row of Figure 10 presents the generated images. The images derived from both classes have graffiti styles, which means that the model passes the face validity check. Moreover, it is apparent that the generated images differ in style across the two classes, with the least-preferred class being darker in color, and the most-preferred class being brighter. This pattern is consistent with the pattern observed in the internal images.
Figure 10: CcGAN-Generated Images
CcGAN with Internal Preferences Only
Refer to caption Refer to caption
(a) Least-Preferred Class (b) Most-Preferred Class
CcGAN with Internal and External Preferences
Refer to caption Refer to caption
(a) Least-Preferred Class (b) Most-Preferred Class
Similarly, we present the new designs generated for the least-preferred class and the most-preferred class by the CcGAN model with both internal and external preferences. As shown in the second row of Figure 10, the generated images differ in style across the two classes. The difference is more salient than in the case in which only internal preferences are incorporated. The least-preferred class is much darker in color, and the most-preferred class is brighter and has more blue/reddish colors. The image patterns in the most-preferred class also more exemplify the graffiti theme than do those in the least-preferred class. This means that the external preferences help generate images with higher quality in terms of mimicking the graffiti features in the most-preferred class.
As our CcGAN models utilize continuous popularity labels, they are able to generate a range of images with continuous popularity levels. Figure 11 displays the progression of the generated images ranging from the lowest popularity level to the highest popularity level. The top row shows the images generated solely with the internal preferences, while the bottom row shows the images generated with both internal and external preferences. A noticeable trend is the gradual change in the style of the generated images as the popularity level increases in both models. Images with higher popularity levels have brighter colors and clearer patterns, which resemble the observed patterns among the real images.
Figure 11: Generated Graffiti Images with Continuous Labels
Note: The top row shows the images generated by the CcGAN with internal preferences only, while the bottom row shows the images generated by the CcGAN with both internal and external preferences.
In summary, the DCGAN, the CcGAN with internal preferences, and the CcGAN with both internal and external preferences all generate meaningful graffiti templates that have similar styles to those of the internal templates. Furthermore, the CcGAN models are able to generate new templates with different popularity levels. This ability is stronger for the CcGAN with both internal and external preferences, highlighting the benefits of incorporating external preferences into the generation task.
5.3 Performance of the Generative Models
After establishing the validity of the generative models in terms of their ability to generate new designs in a controlled manner, we now compare their performances with respect to generating new appealing designs that are likely to be preferred by consumers, which is the main goal of the proposed framework.
Performance based on model-predicted popularity
Given the newly generated designs, we assess how attractive they are to consumers and how likely consumers are to choose them. Our prediction model is trained to predict consumer choices, so we can use our trained prediction model to calculate consumers’ choice probabilities for the new designs generated by the three generative models.
Table 3 shows the average predicted choice probabilities of the internal real templates, DCGAN-generated templates, and CcGAN-generated templates (from the most preferred class). The comparison illustrates whether the generative models can generate new designs that outperform the existing offerings of the company. It also illustrates which generative model can generate the most appealing new designs. As we have two prediction models, we use Prediction Model 1 to evaluate the attractiveness of the templates in the first row and use Prediction Model 2 in the second row. To maintain consistency between the prediction and generation models, we use Prediction Model 1 (Prediction Model 2) to evaluate the performance of the CcGAN with internal preferences only (with internal and external preferences). To facilitate the comparison, we normalize the choice probabilities of the real internal images to 1 in the first column and present the percentage changes relative to the first column in the other columns.
The results show that the new DCGAN-generated designs do not perform well when evaluated by Prediction Model 1. Intuitively, the traditional GAN method (i.e., the DCGAN) only incorporates external templates when generating new templates. Although the external templates help bring new design features to the DCGAN-generated templates, Prediction Model 1 cannot recognize the attractiveness of such new features because it learns only consumer preferences on internal templates. The external features that are different from the popular internal features are treated as unpopular by Prediction Model 1. Therefore, the performance of the DCGAN model is even worse than that attained on the internal real images (-4.43%).
The CcGAN model with internal preferences exhibits better performance by using the internal preferences to guide the generative model, allowing it to generate templates that are more appealing to internal consumers. However, this improvement is limited (0.72%) because Prediction Model 1 is trained only on internal preferences and cannot fully recognize the out-of-the-box features derived from the external templates.
Further incorporating external preference information allows the CcGAN to fully leverage the appealing features of the external templates and substantially improve the performance of the generative model. The incorporation of external preference information allows Prediction Model 2 to learn the types of features that consumers prefer in the external templates. The underlying assumption is that external consumers with similar appearances to those of the internal consumers have similar preferences, so the preference information of the external consumers can be used to infer the preferences of the internal consumers. Prediction Model 2 is thus able to tell which external features might be appealing to the internal consumers, guiding the CcGAN to incorporate popular features acquired from both internal and external sources and leading to the largest model performance improvement (5.37%). The performance of the DCGAN is also better (3.54%) when evaluated by Prediction Model 2 because this model can now recognize popular features derived from external sources in the DCGAN-generated templates.
Table 3: Average Predicted Choice Probabilities
Evaluation Methods Internal DCGAN CcGAN CcGAN
Real Images (Internal) (Internal and External)
Prediction Model 1 1.00 -4.43% 0.72% N.A.
Prediction Model 2 1.00 3.54% N.A. 5.37%
Note: This table presents the average choice probabilities predicted for real images and images generated using different GAN models. The first (second) row shows the choice probabilities predicted by Prediction Model 1 (Prediction Model 2). We normalize the first column to 1 (i.e., make the choice probability of the real image equal 1) and present the other columns as percentage improvements relative to the first column.
In addition to the average predicted choice probabilities in Table 3, we also present the histograms of the predicted choice probabilities for the internal real images, DCGAN-generated images, and CcGAN-generated images in Figure 12. Specifically, Figure 12 (a) shows the choice probabilities predicted by Prediction Model 1, and Figure 12 (b) shows the choice probabilities predicted by Prediction Model 2. The findings from the histograms further confirm the findings in Table 3. First, incorporating consumer preference information is helpful because the CcGAN always outperforms the DCGAN. Second, incorporating external consumer preferences is particularly helpful. The CcGAN-generated images are not significantly better than the real internal images if only internal preferences are incorporated. However, when both internal and external preferences are added, the CcGAN-generated images are significantly better than the real internal images.
Overall, the results highlight the importance of incorporating internal and external preference information. The external preference information is particularly powerful. It allows the predictor model to fully recognize the external features that are potentially appealing to the internal consumers, thereby guiding the generation process to fully incorporate these attractive external features.
Figure 12: Histograms of the Predicted Choice Probabilities
Refer to caption Refer to caption
(a) Predicted by Prediction Model 1 (b) Predicted by Prediction Model 2
Note: This figure plots the histograms of the predicted average choice probabilities for real images and images generated using different GAN models.
Performance based on outside metrics
The previous performance test used a prediction model within our framework to show the popularity of the generated images. We further use metrics generated outside our framework to evaluate the popularity of the generated images. Metrics outside our framework can serve as more objective and independent measures, providing further support for our results.
We propose a distance-based metric as an indicator of popularity. Intuitively, popular internal templates have common features, suggesting that consumers may favor certain types of templates. Therefore, we can look at how similar (far way) the generated images are to the popular (unpopular) internal templates. The more closely the newly generated templates resemble the known popular templates, the more likely that they are also appealing to consumers.
The calculation procedure of the distance-based metric is as follows. First, we identify the three most-preferred internal templates $\mathcal{J}=\{j_{1},j_{2},j_{3}\}$ and the three least-preferred templates $\mathcal{K}=\{k_{1},k_{2},k_{3}\}$ from the observed consumer choices. Second, we calculate a 1000-dimensional vector embedding for the six identified internal templates and the new templates generated by different generative models. Third, we calculate the distance between each generated template $i$ and the most- and least-preferred internal templates. We further normalize the distances to be within the range of [0, 1]. Let $d_{ij}$ and $d_{ik}$ denote the normalized distances between template $i$ and the most-preferred template $j$ and the least-preferred template $k$, respectively. Finally, we calculate the distance-based metric for each generated template and average the results over all templates generated from the same generative model to obtain a measure for each generative model. To summarize, the distance-based metric for each generative model is defined as
$D=\frac{1}{N}\sum_{i=1}^{N}\Big{\{}\sum_{j\in\mathcal{J}}d_{ij}^{2}+\sum_{k\in\mathcal{K}}(1-d_{ik})^{2}\Big{\}},$
where $N$ denotes the total number of templates generated by each generative model. This distance-based metric decreases when the generated templates are more (less) similar to the most-preferred (least-preferred) internal templates. The smaller the distance-based metric is, the better. The metric provides an independent assessment of model performance because it does not rely on the prediction model within our framework. Instead, it leverages the intrinsic similarity among the image features to suggest popularity.
Table 4 presents the distance-based metric results produced by different generative models. We find that the distance metric is smaller for the CcGAN with internal preferences and smallest for the CcGAN with both internal and external preferences. The results suggest the value of incorporating internal and external preferences into the process of generating new templates. Figure 13 provides histograms for the distance-based metrics yielded by different generative models. This result further confirms that the CcGAN with both internal and external preferences performs best.
Table 4: Distance-Based Metrics Produced for the Generated Images
Statistics DCGAN CcGAN (Internal) CcGAN (Internal and External)
Mean 0.58 0.54 0.44
Median 0.55 0.52 0.40
Note: This table presents the mean and median values of the distance-based metrics produced for images generated using different methods. A small value means that the generated templates are close to the popular internal templates while far away from the unpopular internal templates in terms of visual characteristics. A small metric suggests a higher degree of popularity among consumers.
Figure 13: Histograms of the Distance-Based Metrics
Note: This figure plots the histograms of the distance-based metrics for the images generated using different GAN models. A small value means that the generated templates are close to the popular internal templates while far away from the unpopular internal templates in terms of visual characteristics. A small metric suggests a higher degree of popularity among consumers.
6 Managerial Implications and Conclusion
In this paper, we provide a practical solution for the product design decisions of firms using state-of-the-art deep learning methods (i.e., GANs). The unique contribution of our work is that our proposed model systematically incorporates consumer preference information and external UGC information into the design generation process. First, we construct popularity labels based on a prediction model trained on historical consumer choices and include the labels as additional inputs to guide the training process of the generative model. Such an approach enables the generative model to directionally generate new product designs that are more likely to be preferred by consumers. Second, our model allows firms to leverage the rich design ideas and consumer tastes contained in UGC. Both the predictor model and the generative model are trained on external product designs and external consumer preferences. The external images not only serve as extra training resources for the generative models but also provide additional external consumer preference information that can be utilized to guide the generative model.
Two important managerial takeaways can be derived from our results. First, it is beneficial to incorporate consumer preference information into the generation process because it can significantly improve the popularity and quality of the generated designs. Second, it is not sufficient to only incorporate external designs; external preference information is also required for the generative model to fully incorporate new appealing features from external designs. When the predictor model is trained only on internal preferences, consumers cannot understand how to evaluate the new features contained in the external designs. Therefore, the predictor model cannot guide the generator model to fully incorporate new external features. Training the predictor model with both internal and external preferences can fully release its power, enabling it to guide the generative model. By doing so, industry practitioners can better capture consumers’ heterogeneous preferences for various styles of product designs, especially those that are not present in the existing internal designs.
Our proposed model offers a superior alternative to the prevailing industry practice that relies solely on pure GAN models without considering consumer preferences. The proposed model is cost-effective, automated, and scalable. It also generates new designs in a controllable and targeted way. The new designs are more likely to be preferred by consumers ex ante, which contrasts with the approach of generating a larger set of new designs and screening unpopular designs on an ex post basis. In this sense, our approach avoids wasting time and resources on new designs that are later shown to be less appealing.
Our model provides an effective solution to the product design ideation problem. The traditional product design ideation method relies on professionally trained human designers, who have limited capacity and may not be able to generate high-quality designs on a large scale. In addition, human designers may not always understand the most recent consumer requirements and cannot systematically learn the preferences of many consumers. We show that the existing designs acquired from external sources can complement the traditional human design process. UGC images can be regarded as design ideas created by individual users in real time and serve as a rich ideation source.
The proposed model offers a cost-effective alternative for obtaining consumer evaluations of new product designs. In contrast to traditional methods such as conducting consumer surveys, we leverage UGC as a source of consumer preferences. This UGC serves as a vast repository of real-time, self-evident consumer preferences that can be utilized to inform design decisions. Such information is highly relevant because it is voluntarily generated by consumers. This information is abundant, readily available online, and dynamically updated, which allows firms to keep up with constantly evolving consumer demands. UGC is especially useful for start-up firms that do not have sufficient internal resources and cannot afford the costs of designing new products using traditional methods. The use of UGC not only reduces the financial burden associated with extensive consumer surveys but also provides a more dynamic and authentic reflection of consumer tastes.
In summary, our findings suggest that industry practitioners can benefit significantly from combining preference information with generative models, moving beyond the limitations of pure GAN-based approaches. By recognizing the importance of external preferences, particularly those embedded in UGC, businesses can enhance the efficiency of their design processes and gain valuable insights into consumer preferences without conducting costly surveys.
Our method has broad applications not only for artistic designs but also for other design tasks where heterogeneous consumer preferences widely exist. Future works can study how consumer preferences can be incorporated into generative models in other contexts and evaluate the effectiveness of such approaches.
Appendix A Additional Tables and Figures
Figure A.1: Illustration from the OpenFace Project
Figure A.2: Illustration of the Reference BVLC Caffenet
Figure A.3: Histograms of the Predicted Popularity Levels (Internal Preferences)
Refer to caption Refer to caption
(a) Original Popularity Level (b) Grouped Popularity Labels
Figure A.4: Histograms of the Predicted Popularity Levels (Internal and External Preferences)
Refer to caption Refer to caption
(a) Original Popularity Level (b) Grouped Popularity Labels
Appendix B Analysis of the Traditional Chinese Theme
To further validate our results, we show that our proposed model works on a wide range of themes by applying it to generate templates belonging to the traditional Chinese theme. Similar to the graffiti case, we generate new traditional Chinese theme-based designs using the baseline DCGAN model without consumer preferences, the enhanced model where we adopt the CcGAN with internal consumer preferences, and the advanced model where we adopt the CcGAN with internal and external consumer preferences.
Similar to the graffiti theme, we gather photos from social media and image websites featuring human subjects with styles matching the internal templates of the traditional Chinese theme. These external images are then preprocessed by removing the human subjects and completing the backgrounds through inpainting, ensuring that they match the style of the internal templates offered by the company. Simultaneously, we retain the removed faces for potential later use. This screening and inpainting process yields a total of 110 traditional Chinese images. Figure B.5 provides examples of both internal and external photos that are suitable for the traditional Chinese theme. As depicted, the collected external images exhibit a similar style to that of the internal images and showcase adequate variations.
Refer to caption Refer to caption
(a) Internal images (b) External images
Figure B.5: Examples of Traditional Chinese Templates: Internal vs. External
B.1 Performance of the Prediction Model
As discussed in Section 4.1, during the training process of the prediction model on internal data, all the themes within the “main (single)” category, including the traditional Chinese theme, are utilized. Hence, there is no need to retrain Prediction Model 1, as the results obtained for both the graffiti and traditional Chinese themes are derived from the same model. The focus is on retraining Prediction Model 2, which incorporates 110 external photos as positive samples. With the addition of these external images to the 33,804 internal images, the sample size totals 33,914. Following the previous approach, 70% of the observations are allocated to the training set. To mitigate the unbalanced sample issue, the minority class is upsampled to 20,000 samples, aligning with the majority class. Prediction Model 2 is estimated using a random forest model with the same hyperparameters as those employed in the graffiti case. Table B.1 displays the accuracy of the prediction model incorporating both internal and external preferences. The results indicate that the inclusion of external preferences has a limited impact on the overall accuracy, primarily due to the restricted size of the external traditional Chinese image dataset.
Table B.1: Accuracies Achieved by the Consumer Choice Model Prediction On the Test Set
Choice Prediction Task Acc. Balanced Acc. FNR FPR
Internal Preferences Only 0.793 0.791 0.212 0.206
Internal and External Preferences 0.794 0.790 0.216 0.205
Note: This table presents the accuracy, balanced accuracy rate, false-negative rate, and false-positive rate produced for the consumer choice model prediction results on the test set. The first row presents the accuracy for the prediction model with internal preferences only, and the second row presents the accuracy for the prediction model with both internal and external preferences, where the external preferences are derived from the external traditional Chinese theme.
We then conduct a preliminary visual assessment to evaluate the effectiveness of Prediction Model 1 in terms of generating meaningful popularity labels. To do so, we identify the most- and least-preferred internal templates based on the consumer adoption rates. We then compare their styles with those of the combined internal and external templates that have the highest and lowest average predicted choice probabilities, respectively. As illustrated in Figure B.6, we observe that the external templates with high choice probabilities closely resemble the most popular internal templates. Conversely, the external templates with low choice probabilities are similar to the least-popular internal templates. This observation visually validates the accuracy of the predicted popularity labels.
Figure B.6: Comparing the Internal and External Templates by Popularity (Internal Preferences)
Most-Preferred Templates Least-Preferred Templates
Internal Templates Refer to caption Refer to caption
Combined Templates Refer to caption Refer to caption
We also conduct a preliminary visual assessment to evaluate the performance of Prediction Model 2 in terms of generating meaningful popularity labels when integrating both internal and external preferences. By comparing the most- and least-preferred internal templates with their counterparts predicted by the model (Figure B.7), we observe that the external templates with high choice probabilities exhibit styles similar to those of the most-popular internal templates, while those with low probabilities resemble the least-popular internal templates. However, in contrast to Figure 8, the similarity is less pronounced, as the new prediction model captures some external consumer preferences that are not observed in the internal data, particularly a preference for backgrounds featuring blossoms and birds. Consequently, the images with these elements are detected through the incorporation of the external preferences. Nonetheless, certain images remain consistent across both models; for example, the third images of the most-preferred external template are identical in both cases. Similarly, the third images of the least-popular external template are also consistent across models. This pattern confirms that the enhanced prediction model leverages new information derived from external preferences, enabling it to better capture consumer preferences for producing more diverse product designs.
Figure B.7: Comparing the Internal and External Templates by Popularity (Internal and External Preferences Together)
Most-Preferred Templates Least-Preferred Templates
Internal Templates Refer to caption Refer to caption
Combined Templates Refer to caption Refer to caption
B.2 Performance of the Generative Models
We first train the DCGAN model by combining the existing templates derived from internal and external sources, resulting in 119 training images for the traditional Chinese theme, 9 of which are acquired from internal data and 110 of which are acquired from external data. Since the size of our training dataset is extremely small, we again adopt the DiffAugment method (Zhao et al., (2020)), following the approach used in the graffiti case. All other hyperparameters remain consistent with those employed in the graffiti case.
Next, we proceed with the enhanced model, training the CcGAN with internal consumer preferences. Following our previous approach, we group images with similar popularity levels into bins, ensuring that the images within the same bin share an identical popularity level. Any class with fewer than 50 images is replicated to ensure a minimum of 50 images per class, thereby achieving a more balanced distribution. Subsequently, we normalize each label by dividing it by the maximum label value, resulting in 7 normalized labels ranging from 0.586 to 1. We employ the soft vicinal discriminator loss (SVDL) as the discriminator loss, with $\sigma$ and $\kappa$ selected using the rule-of-thumb formulae. The CcGAN undergoes training for 20,000 iterations with a constant learning rate of $10^{-4}$ and a batch size of 64. The latent GAN dimensionality is set at 256, and the discriminator updates twice per iteration. To address the small sample size problem, we utilize the DiffAugment method during training, as previously done in the DCGAN case. Once the CcGAN is trained, we use the generator to produce new template designs with various popularity labels, where label=0.586 represents the least-preferred class and label=1 indicates the most-preferred class.
Finally, we train the advanced model: the CcGAN incorporating both internal and external consumer preferences. Similar to the enhanced model, we utilize Prediction Model 2 to determine popularity levels for the existing templates. Continuous labels are grouped into 5 bins, with any groups containing fewer than 20 images replicated to ensure a minimum of 20 images per group. The labels are then normalized, resulting in 5 values ranging from 0.529 to 1. The SVDL is used again with the hyperparameters obtained by following the previous guidelines. After training, the generator produces new template designs with various popularity labels, ranging from least preferred (label=0.529) to most preferred (label=1).
Images generated using different models
We start by presenting the performance of the DCGAN models. Figure 8(b) compares the images generated by the trained DCGAN models from the real images belonging to the traditional Chinese theme. Figure 8(b) (a) shows the real traditional Chinese images, where these templates feature a light-colored background with some plum blossoms. As shown in Figure 8(b) (b), the styles of the DCGAN-generated images are similar to those of the real images. Significant variations are also observed in the styles of these generated images. Therefore, by employing a simple visual examination, we know that the DCGAN model does a good job when generating traditional Chinese theme templates.
Figure B.8: Real vs. DCGAN-Generated Images Belonging to the Traditional Chinese Theme
(a) Real Images
(b) DCGAN-Generated Images
Next, we show that the CcGAN models can controllably generate new designs with different popularity levels. Figure B.9 shows the generated images, ranging from the least-preferred class to the most-preferred class. The top row displays the images generated using internal preferences in the CcGAN models, while the bottom row shows the images generated with both internal and external preferences. A clear trend emerges: as the label increases, the style of the generated images gradually shifts, indicating alignment with the consumer preferences across different classes.
Figure B.9: Traditional Chinese Images Generated with Continuous Labels
B.3 Comparing Different GAN Models
We now compare the results of these three methods and check whether any positive effect results from incorporating consumer preferences into the process of generating new designs. Table B.2 presents the choice probabilities predicted for the generated images using various methods. The first column is normalized to 1, representing the choice probability of the real images, while the subsequent columns display their percentage improvements relative to the first column. The first row illustrates the choice probabilities predicted when only internal preferences are incorporated into the prediction model, while the second row depicts the choice probabilities predicted when both internal and external preferences are considered. The table compares the DCGAN, the CcGAN with internal preferences, and the CcGAN with both internal and external preferences. The DCGAN relies solely on external templates, resulting in a decrease in performance (-4.55%) compared to that achieved with internal real images. This decline is attributed to the inability of the predictor model to recognize potentially appealing features from external templates. The CcGAN with internal preferences yields improved performance by guiding a GAN with internal preferences, yet its efficacy is constrained because the predictor model fails to train on external features. Incorporating external preference information enhances the performance of a GAN by leveraging the appealing features acquired from external templates. The predictor model learns the preferred features from external consumers, assuming that the preferences of internal and external consumers are similar. Consequently, the predictor-guided CcGAN provides the largest improvement (11.35%), generating templates with popular features from both sources. The performance of the DCGAN also improves (8.97%) as the predictor model recognizes popular external features in the generated templates.
Table B.2: Predicted Choice Probabilities for the Generated Images
Methods Internal Real Images DCGAN CcGAN (Label=1)
Internal Preferences Only 1.00 -4.55% -0.84%
Internal + External Preferences 1.00 8.97% 11.35%
Note: This table presents the average selection probabilities predicted for the images generated using different methods. We normalize the first column to 1 (i.e., make the selection probability for real images=1) and present the other columns as percentage improvements relative to the first column. The first row calculates the selection probabilities for the internal real images, DCGAN-generated images, and CcGAN-generated images (with internal preferences incorporated) for the most-preferred class (label=1) using the prediction model where only internal preferences are added. The second row calculates the selection probabilities for the internal real images, DCGAN-generated images, and CcGAN-generated images (with internal and external preferences incorporated) for the most-preferred class (label=1) using the prediction model with both internal and external preferences added.